EARTH’S MODEL WITH VARIABLE CHANDLER’S FREQUENCU

Nadežda Pejović

Astronomical institute, Faculty of Sciences, University of Belgrade, Studentski trg 16, 11000 Beograd, Yugoslavia

The hypothesis of a nonlinear dependence of Chandler’s frequency on the total polar motion amplitude is confirmed by analysing the ILS coordinates of the pole. Assuming that this nonlinear dependence of Chandler’s frequency is due to a nonequilibrium ocean, one introduces a new excitation function of the nonequilibrium ocean in Vondràk’s (1987) model of the Earth. The responce of this more complete Earth’s model with a variable and amplitude-dependent Chandler’s frequency to

atmospheric excitation is derived, as well.

Key words: Chandler’s frequency – polar motion – atmospheric excitation